OpenStudy (anonymous):

A color printer prints 15 pages in 7 minutes. How many minutes does it take per page? If necessary, round your answers to the nearest hundredth.

OpenStudy (anonymous):

\[\frac{ 15 \ pages }{ 7 \ minutes } = \frac{ 1 \ page }{ \frac{ 7 }{ 15 } \ minutes}\ = \frac{ 1 \ page}{ .47 \ minutes }\]